OpenStudy (ria23):

y = -x2 + 8x - 12 a.) Identify the location of the vertex b.) Identify the location of the x-intercept(s), if any c.) Identify the location of the y-intercept d.) Write the equation for the axis of symmetry (I'm having trouble converting this from standard to the other forms.)

TheSmartOne (thesmartone):

Do you know how to find the vertex? Do you know what an x and y intercept is?

OpenStudy (ria23):

I thought I had the formula for vertex, but I guess I had the formula for point slope and was getting them mixed up... >.< I don't know the formula for vertex

TheSmartOne (thesmartone):

Vertex=\[(\frac{ -b }{ 2a},f(\frac{ -b }{ 2a}))\] in the form of (x,y)

OpenStudy (ria23):

I just plug in a and b and find it and that's my vertex?

TheSmartOne (thesmartone):

yes and then whatever you get from -b/2a you plug that back in to find the y value.

TheSmartOne (thesmartone):

Or you can graph it and then see where everything is...

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/vvrjzcunih

TheSmartOne (thesmartone):

You will notice that the vertex is (4,4)

OpenStudy (ria23):

Would it be 4 or -4?

OpenStudy (ria23):

Oh :D haha...

TheSmartOne (thesmartone):

Look at the graph.. You can see the vertex and even the intercepts...

OpenStudy (ria23):

Yes. c: How would I calculate the intercepts?

TheSmartOne (thesmartone):

y intercept is when x=0 and you solve for y x intercepts is when y=0 and you solve for x

TheSmartOne (thesmartone):

but if you look at the graph you see the x-intercepts is (2,0) and (6,0) and the y-intercepts is (0,-12)

OpenStudy (ria23):

So, I would set x=0 ... To find the y intercept? Or the x?

TheSmartOne (thesmartone):

Axis of symmetry is \[x=\frac{-b}{ 2a }\]

TheSmartOne (thesmartone):

Yes you would make x=0 to find the y-intercept and solve for y and your point would be (0,y)

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/mrlwubqoo8

TheSmartOne (thesmartone):

This graph includes the axis of symmetry.

TheSmartOne (thesmartone):

So do you have any more questions?

TheSmartOne (thesmartone):

@Ria23

OpenStudy (ria23):

Uhm... I think, I have it... If I get stuck anywhere I'll tag yhu c: Thank yhu for helping me! I appreciate it!

TheSmartOne (thesmartone):

no problem...\[\huge\color{red}{\ddot\smile}\]

TheSmartOne (thesmartone):

@ria23 If your question is solved.. you can close this question...

OpenStudy (ria23):

@TheSmartOne I'm having trouble identifying how I find the x-intercepts. If I set y=0, that'll give me 0=x^2+8x-12 How would I solve for the x's... I have my graph drawn, but I need to show my steps on how I got it